# EmojiDiff: Advanced Facial Expression Control with High Identity Preservation in Portrait Generation

### v1

@@ 3.5 ID-enhanced Contrast Alignment
Both identity control and expression control manipulate the face of the generated image, causing potential interference between the two. Inevitably, incorporating expression control affects the original identity control in portrait generation tasks. Our findings indicate that emphasizing expression control alters muscle details and facial attributes, such as the presence of glasses, due to the generative diversity inherent in the diffusion model.
To this end, we propose ICA to fine-tune the E-Adapter. Based on the same inputs, ICA incorporates the frozen Refined E-Adapter and the original identity controller (IP-Adapter [47]) as supervision to calculate expression loss $\mathcal{L}_{exp}$ and identity loss $\mathcal{L}_{id}$. $\mathcal{L}_{id}$ encourages the E-Adapter which integrates expression control, and the IP-Adapter to generate similar portraits, thereby minimizing the impact of expression control on identity fidelity. Additionally, $\mathcal{L}_{exp}$ ensures that fine-tuning does not compromise the existing expression control capability of the Refined E-Adapter.
Both losses rely on the denoised clean image $\hat{x}_{0}$. Contrary to standard diffusion inference which obtains clean images in multiple steps, we achieve approximately reconstructed images through one-step decoding. Specifically, we propose the Adaptive Noise Inversion (ANI) to directly acquire $\hat{x}_{0}$ at any given timestep $t$, denoted by the formula:
@@
where ”${\rm Decode}$” is the vae decode function, and $R_{tmax}$ is a constant. When $t>R_{tmax}$, we set $f(t)$ to $f(R_{tmax})$ to prevent reconstructed sample $\hat{x}_{0}$ from becoming too noisy, which compromises the accurate calculation of losses. Derivation details can be found in supplementary materia.
Building on ANI, we employ the IP-Adapter to generate portrait image $\hat{x}_{0}^{id}$, and the currently fine-tuning E-Adapter to generate the image with expressions $\hat{x}_{0}^{cur}$. $\mathcal{L}_{id}$ minimizes the identity difference between two images, is defined as:
@@
Table 1: Quantitative comparisons of our method with SOTA counterparts in various base models.
@@ 4 Experiments
We first give an overview of the dataset, implementation details, baselines, and benchmarks used in the experiments. Then, we present the experimental results on different base models, followed by an ablation study of each module.
@@
Dataset. We first collect 10k indoor and outdoor images of people of different countries and genders, each with 3-12 expressions. Then, we use LIQE [50] to filter out low-quality images and crop out 512$\times$512 face square images. Subsequently, this dataset is employed to train the Base E-Adapter based on SD1.5 [37]. Afterward, we construct about 120k cross-identity triplets through IDI. Through expression consistency filtering, we retain about 100k triplets for training and fine-tuning the Refined E-Adapter. For evaluation, we gather images of 50 celebrities from diverse fields, with 30 different expressions.
@@
Implementation Details. We train the Base E-Adapter on SD1.5 to construct data and the Refined E-Adapter on both SD1.5 [37] and SDXL [35]. In the Base E-Adapter training stage, our model is trained from scratch for 0.5 days. In SD1.5, we train the Refined E-Adapter for 1 day and fine-tune it for 4 hours. In SDXL, the input images are super-resolution using SwinIR [26] to a resolution of 1024$\times$1024, and we train the Refined E-Adapter for 2 days and fine-tun it for 0.5 days. We select Up 0.1. as the expression layers in SDXL. We do not observe style leakage on SD1.5, so we choose all its layers as expression layers.
@@
Baselines. We compare our model with prior methods including OpenPose [49], MediaPipe [29], Arc2Face [32], and FineFace [42], using the same configurations (base model, text prompt, inference steps, etc.). Although portrait animation methods are fundamentally different from our approach, we assess these methods, including FaceAdapter [18], X-Portrait [45], and LivePortrait [16], to thoroughly evaluate our method. For the diffusion-based methods, we integrate IP-Adapter [47] plugin for identity preservation. For animation-based methods, an expressionless image of the portrait is generated, and then animation is applied. We select LivePortrait [16] to construct data triples $\mathcal{T}$ and train the E-Adapter, which serves as our strong baseline.
@@
Benchmarks. To ensure a fair comparison, all experiments are conducted using the SD1.5 framework across three styles: realistic, anime, and ink painting. Given the identity image, expression reference, and generated image, we evaluate all methods using ID fidelity (ID), image quality (IQ), expression similarity (Exp.), and facial landmarks movements similarity (LMS). Different from the models used in the ID encoder, the Antelopev2 [1] model is utilized to extract identity embeddings from both the portrait and generated image, with the cosine similarity between them used to evaluate identity fidelity. For the assessment of image quality, the pre-trained network LIQE [50] is employed. Expression similarity is measured by computing the L1 difference between facial blendshapes extracted from the expression and generated image using MediaPipe [29]. To evaluate the keypoint similarity, crucial facial landmarks (such as eyes, pupils, and mouth) are detected with MediaPipe to calculate movement amplitude differences. More metric details are provided in supplementary materia.
@@ 4.1 Comparison
Qualitative results. Fig. 6 illustrates the performance of various methods in ID and expression control across three distinct styles. These examples demonstrate that our model effectively conveys reference expressions, including lip movements and eye gaze while preserving the identity information of the source portrait. The fourth and sixth examples in the figure highlight the ability of our method to capture subtle movements, including lower lip pursing and pouting. Compared to other approaches, our model consistently exhibits robust expression control across different styles.
@@ 4.2 Ablation Study
We evaluate the proposed IDI and ICA on SD1.5 and SDXL. As presented in Tab. 2, in comparison to our baseline, IDI substantially improves expression control, reducing the Exp. score of the two base models by 23.0% and 22.6%, respectively. Additionally, ICA further enhances ID fidelity without compromising expression control, increasing the corresponding ID score by 3% and 6%, respectively.
@@
ID-irrelevant Data Iteration. As shown in Tab. 3, we assess the performance of E-Adapter at different data iteration stages. Initially, the Base E-Adapter accurately transfers expressions to generated images compared to the baseline, achieving significant reductions in Exp. and LMS scores. Additionally, using the Base E-Adapter for cross-identity data construction, the newly trained Refined E-Adapter substantially reduces ID leakage, improving the ID score by 0.136. Finally, through further face-swapping, the Refined E-Adapter significantly outperforms the baseline.
Expression Controller Architecture. As illustrated in Fig. 7, we experiment with various expression controller structures. Expression embeddings are injected into the diffusion model using one of the following strategies: (a) self-attention, (b) ControlNet, (c) serial cross-attention with ID embeddings, and (d) parallel cross-attention with ID embeddings. As indicated in Tab. 4, the schemes (a) and (c), which operate without synchronous integration with ID embeddings, are insufficient in both ID fidelity and expression control capability. While the ControlNet achieves high ID fidelity, it lacks satisfactory expression control. Our approach combines identity and expression embeddings within a unified cross-attention block, effectively fostering their interaction and delivering optimal performance.
Expression Layer. As illustrated in Fig. 8, applying expression embeddings across all layers in highly stylized models significantly deteriorates the quality, and restricting expression feature injection to only the face layers does not improve it. By confining the injection of expression embeddings to the expression layers, we effectively address this issue. Notably, compared to employing expression embeddings across all layers or solely on face layers, our approach does not impair the expression control capability.
@@
Figure 8: Ablation study of expression layers. Unlike our method, integrating expression information into either the face layers or the entire U-Net layers results in a degeneration in both image style and quality.
@@ 5 Conclusion
We introduce EmojiDiff, an innovative solution that utilizes both portrait images and RGB expression images to generate expressive portraits. Initially, we propose ID-irrelevant Data Iteration to synthesize high-quality cross-identity data pairs for training, effectively decoupling expression and identity information in expression images. Additionally, style leakage from expression signals is prevented by injecting expression image features exclusively into expression-sensitive layers. To further harmonize identity and expression control, we propose ID-enhanced Contrast Alignment to mitigate the impact of expression modulation on identity fidelity. Experimental results demonstrate that our approach achieves accurate and nuanced facial expression control while maintaining precise identity resemblance.
@@ B.1 Detailed Implementation
We gather images of over 10,000 people displaying various expressions from video clips and in-house face databases, removing low-quality images using LIQE [50]. Then, we employ the proposed IDI to construct cross-identity, same-expression datasets. As mentioned in Sec. 3.3 of the main body, facial blendshapes [15] and landmark [29] differences are utilized to filter out expression-changed data during data construction. Specifically, we calculate the Exp. and LMS metrics of the original expression image and the synthesized new image, excluding the data with Exp. $\leq$0.05 and LMS $\leq$0.18. Based on synthesized data, we train the Refined E-Adapter for SD1.5 at a resolution of 512$\times$512. On SDXL, we upscale the data to a resolution of 1024$\times$1024 and train it with random scaling following IP-Adapter [47]. The learning rates of the Adam optimizer during training and fine-tuning are set to 2$\times$ $10^{-5}$ and 5$\times$$10^{-6}$, respectively, with $\beta_{1}$ = 0.5, and $\beta_{2}$ = 0.999.
@@ C Additional Experimental Results
C.1 Attention Map Visualization
As shown in Fig. 1, we present additional visualizations of attention maps. Consistent with Fig. 4 in the main text, the face layers exhibit a clear response in the facial area. At the same time, the expression layers (e.g., Up 0.1.0, Up 0.1.4) focus on various facial regions, enabling comprehensive expression control across all face areas.
C.2 More Visualization Results
In Fig. 6 of the main body, we only display a few generated images due to the page limit. In this subsection, we present additional results generated by our method on more styles and expressions. As shown in Fig. 2 and 3, our method accurately transfers the reference expression to the generated image while maintaining high identity fidelity.
C.3 Data Iteration Visualization
@@
Figure 4: More examples of Stage II. ID-irrelevant Data Iteration (IDI) is introduced to transform expression reference images into the restruct and faceswap images, thus synthesizing high-quality data pairs with differing identities and consistent expressions.

### v2

@@ 3.3 ID-enhanced Contrast Alignment
Both identity control and expression control manipulate the face of the generated image, causing potential interference between the two. Inevitably, incorporating expression control affects the original identity control in portrait generation tasks. Our findings indicate that emphasizing expression control alters muscle details and facial attributes (e.g., the presence of glasses, hair, and facial textures), due to the generative diversity inherent in the diffusion model.
To this end, we propose ICA to fine-tune the E-Adapter. Based on the same inputs, ICA incorporates the frozen Refined E-Adapter and the original identity controller (IP-Adapter [49]) as supervision to calculate expression loss $\mathcal{L}_{exp}$ and identity loss $\mathcal{L}_{id}$. $\mathcal{L}_{id}$ encourages the IP-Adapter and the E-Adapter which integrates expression control to generate similar portraits, thereby minimizing the impact of expression control on identity fidelity. Additionally, $\mathcal{L}_{exp}$ ensures that fine-tuning does not compromise the existing expression control capability of the Refined E-Adapter.
Both losses are dependent on the denoised clean image $\hat{x}_{0}$. However, standard diffusion inference incurs significant gradient backpropagation time costs and GPU memory usage, and it will increase linearly with the number of timesteps. To overcome this, we propose Adaptive Noise Inversion (ANI), an efficient one-step decoding strategy, denoted by the formula:
@@
where ”${\rm Decode}$” is the VAE decode function, and $R_{tmax}$ is a constant. Through ANI, we can directly reconstruct images $\hat{x}_{0}$ at any given timestep $t$, even with large timesteps characterized by high noise. We present theoretical analysis of this mechanism in supplementary materia Sec. B.2.
Building on ANI, we employ the IP-Adapter to generate portrait image $\hat{x}_{0}^{id}$, and the currently fine-tuning E-Adapter to generate the image with expressions $\hat{x}_{0}^{cur}$. $\mathcal{L}_{id}$ minimizes the identity difference between two images, defined as:
@@
[…]
@@ 4 Experiments
4.1 Experiment Setting
@@
Dataset. We first collect 10k indoor and outdoor images of people of different countries and genders, each with 3-12 expressions. Then, we use LIQE [52] to filter out low-quality images and crop out 512$\times$512 face square images. Subsequently, this dataset is employed to train the Base E-Adapter based on SD1.5 [38]. Afterward, we construct about 120k cross-identity triplets through IDI. Through expression consistency filtering, we retain about 100k triplets named CIEP100K, for training and fine-tuning the Refined E-Adapter. For evaluation, we gather images of 50 celebrities from diverse fields, with 30 different expressions.
@@
Implementation Details. We train the Base E-Adapter on SD1.5 to construct data and the Refined E-Adapter on both SD1.5 [38] and SDXL [36]. In the Base E-Adapter training stage, our model is trained from scratch for 0.5 days. In SD1.5, we train the Refined E-Adapter for 1 day and fine-tune it for 4 hours. In SDXL, the input images are super-resolution using SwinIR [27] to a resolution of 1024$\times$1024, and we train the Refined E-Adapter for 2 days and fine-tune it for 0.5 days. More details are provided in supplementary materia Sec. B.1.
@@
Baselines. We compare our model with prior methods, including OpenPose [51], MediaPipe [30], Arc2Face [33], and FineFace [43], using the same configurations (base model, text prompt, inference steps, etc.). Although portrait animation methods are fundamentally different from our approach, we assess these methods, including FaceAdapter [19], X-Portrait [47], and LivePortrait [17], to thoroughly evaluate our method. For the diffusion-based methods, we integrate IP-Adapter [49] plugin for identity preservation. For animation-based methods, an expressionless image of the portrait is generated, and then animation is applied. We select LivePortrait [17] to construct data triples $\mathcal{T}$ and train the E-Adapter, which serves as our strong baseline.
@@
Benchmarks. To ensure a fair comparison, all experiments are conducted using the SD1.5 framework across three styles: realistic, anime, and ink painting, while the style is directly dependent on the base model and LoRA. Given the identity image, expression reference, and the generated image, we evaluate all methods using ID fidelity (ID), image quality (IQ), expression similarity (Exp.), and facial landmark movements similarity (LMS). Different from the models used in the ID encoder, the Antelopev2 [1] model is utilized to extract identity embeddings from both the portrait and generated image, with the cosine similarity between them used to evaluate identity fidelity. For the assessment of image quality, the pre-trained network LIQE [52] is employed. Expression similarity is measured by computing the L1 difference between facial blendshapes extracted from the expression and generated image using MediaPipe [30]. To evaluate the keypoint similarity, crucial facial landmarks (such as eyes, pupils, and mouth) are detected with MediaPipe to calculate movement amplitude differences. More metric details are provided in supplementary Sec. A.
@@ 4.2 Comparison
Qualitative results. Fig. 5 illustrates the performance of various methods in ID and expression control across three styles. These examples demonstrate that our model effectively conveys reference expressions, including lip movements and eye gaze while preserving the identity information of the source portrait. The fourth and sixth examples in the figure highlight the ability of our method to capture subtle movements, including lower lip pursing and pouting. Compared to other approaches, our model consistently exhibits robust expression control across different styles.
@@ 4.3 Ablation Study
We evaluate the proposed IDI and ICA on SD1.5 and SDXL. As presented in Tab. 2, in comparison to our baseline, IDI substantially improves expression control, reducing the Exp. score of the two base models by 23.0% and 22.6%, respectively. Additionally, ICA further enhances ID fidelity without compromising expression control, increasing the corresponding ID score by 3% and 6%, respectively. Similarly, as illustrated in Fig. 8, the complete method retains more identity details, including glasses and hair color, compared to the method without ICA.
@@
ID-irrelevant Data Iteration. As shown in Tab. 3, we assess the performance of E-Adapter at different data iteration stages. Initially, the Base E-Adapter achieves significant reductions in Exp. and LMS scores compared to the baseline, showing the powerful expression transfer ability of Base E-Adapter. Additionally, by randomizing the ID branch’s input for cross-identity data construction, the newly trained Refined E-Adapter substantially reduces ID leakage and improves the ID score by 0.136. Finally, through further face-swapping, the Refined E-Adapter significantly outperforms the baseline.
Expression Controller Architecture. As illustrated in Fig. 6, we experiment with various expression controller structures. Expression embeddings are injected into the diffusion model using one of the following strategies: (a) self-attention, (b) ControlNet, (c) serial cross-attention with ID embeddings, and (d) parallel cross-attention with ID embeddings. As indicated in Tab. 4, the schemes (a) and (c), which operate without parallel integration with ID embeddings, are insufficient in both ID fidelity and expression control capability. While the ControlNet achieves high ID fidelity, it lacks satisfactory expression control. Our approach combines identity and expression embeddings within a unified cross-attention block, effectively fostering their interaction and delivering optimal performance.
Generalizability of Expression Control. As depicted in Fig. 7, prompts can be customized to meet diverse requirements while expressions are stably governed by the expression template. Notably, an appropriate prompt will further enhance the vividness of expressions, and a conflicting prompt only slightly affects expressions, as demonstrated in the last two columns of Fig. 7. Moreover, the third and fourth columns of Fig. 8 demonstrate that the identity and expression branch can independently perform their respective functions, enhancing the flexibility of facial control.
@@
Figure 7: Images generated using different prompts.
Figure 8: Images generated without using the ID branch, expression branch, and ICA.
@@ 5 Conclusion
We introduce EmojiDiff, an innovative solution that utilizes both portrait images and RGB expression images to generate expressive portraits. To achieve this, we design a two-stage scheme including decoupling training and fine-tuning. In the training stage, we propose ID-irrelevant Data Iteration to synthesize identity-expression decoupling data pairs, enabling the model to accept the RGB expression template without interference from expression-irrelevant features. In the fine-tuning stage, we propose ID-enhanced Contrast Alignment to further harmonize identity and expression control. This approach ensures the training process through Adaptive Noise Inversion and mitigates the impact of expression modulation on identity fidelity. Experimental results demonstrate that our approach achieves accurate and nuanced facial expression control while maintaining precise identity resemblance.
@@ B.1 Detailed Implementation
We gather images of over 10,000 people displaying various expressions from video clips and in-house face databases, removing low-quality images using LIQE [52]. Then, we employ the proposed IDI to construct cross-identity, same-expression datasets. As mentioned in Sec. 3.2.2 of the main body, facial blendshapes [16] and landmark [30] differences are utilized to filter out expression-changed data during data construction. Specifically, we calculate the Exp. and LMS metrics of the original expression image and the synthesized new image, excluding the data with Exp. $\leq$0.05 and LMS $\leq$0.18. Based on synthesized data, we train the Refined E-Adapter for SD1.5 at a resolution of 512$\times$512. On SDXL, we upscale the data to a resolution of 1024$\times$1024 and train it with random scaling following IP-Adapter [49]. The learning rates of the Adam optimizer during training and fine-tuning are set to 2$\times$ $10^{-5}$ and 5$\times$$10^{-6}$, respectively, with $\beta_{1}$ = 0.5, and $\beta_{2}$ = 0.999.
@@ C Additional Experimental Results
C.1 More Visualization Results
In Fig. 6 of the main body, we only display a few generated images due to the page limit. In this subsection, we present additional results generated by our method on SD1.5 [38] and SDXL [36] framework. As shown in Fig. 1 and 2, our method accurately transfers the reference expression to the generated image while maintaining high identity fidelity.
C.2 Data Iteration Visualization
@@
Figure 3: More examples of IDI. ID-irrelevant Data Iteration (IDI) is introduced to transform expression reference images into the restruct and faceswap images, thus synthesizing high-quality data pairs with differing identities and consistent expressions.